OpenStudy (anonymous):

x = 9 - y 2x - y = -3 (elemination)

OpenStudy (anonymous):

Substitute one into the other, this is easiest: \[2(9-y)-y=-3\] Can you simplify it now and solve for 'y'? Once you do that, plug this 'y' value back into the first equation, 9 minus something. You'll have 'x' too then.

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

np!

jhonyy9 (jhonyy9):

x+y=9 2x-y=-3 --------- 3x 0 =6 x=6/3 =2 x=2

jhonyy9 (jhonyy9):

just because you have wrote there by eliminaton method

jhonyy9 (jhonyy9):

so if x=2 and x+y=9 so 2+y=9 so y=7

OpenStudy (anonymous):

thank you sir :)

jhonyy9 (jhonyy9):

yw good luck bye